OpenStudy (anonymous):

The perimeter of a rectangle is 44ft , and the area of the rectangle is 40ft^2 . Find the dimensions of the rectangle.

OpenStudy (anonymous):

Are you sure about the measurements?

OpenStudy (anonymous):

yes

OpenStudy (aivantettet26):

The equation of the perimeter of a rectangle is P = 2L + 2W and the equation for the area of the rectangle is A = LW Equation 1 A = LW 40 = LW L = 40/W this is our 1st equation We will then substitute equation 1 to the equation of the Perimeter P = 2L+2W 44 = 2(40/W) + 2W Multiply both sides by W W(44 = 2(40/W) + 2W)W 44W = 80 + 2W^2 0 = 2W^2 -44W +80 Use Quadratic Formula 0 = (W - 20) (W - 2) We then have 2 values of W W = 20 or W = 2 Substitute a value of W to the equation of a Area A = LW 40/W = L 40/20 = L 2 = L Substitute the other value of W A = LW 40/W = L 40/2 = L 20 = L We therefore conclude that if the W = 20 then the corresponding Length is 2 and when the W = 2 then the corresponding Length is 20